OpenStudy (anonymous):

a = h (a + B)/2, solve for h

OpenStudy (anonymous):

a = h(a+B) / 2 First multiply each side by 2 2a = h(a+B) Then divide both sides by (a+B) (2a)/(a+B) = h Done! Hope this helps